# Pontryagin Neural Operator for Solving General-Sum Differential Games with Parametric State Constraints

### v1

@@ Abstract
The values of two-player general-sum differential games are viscosity solutions to Hamilton-Jacobi-Isaacs (HJI) equations. Value and policy approximations for such games suffer from the curse of dimensionality (CoD). Alleviating CoD through physics-informed neural networks (PINN) encounters convergence issues when value discontinuity is present due to state constraints. On top of these challenges, it is often necessary to learn generalizable values and policies across a parametric space of games, e.g., for game parameter inference when information is incomplete. To address these challenges, we propose in this paper a Pontryagin-mode neural operator that outperforms existing state-of-the-art (SOTA) on safety performance across games with parametric state constraints. Our key contribution is the introduction of a costate loss defined on the discrepancy between forward and backward costate rollouts, which are computationally cheap. We show that the discontinuity of costate dynamics (in the presence of state constraints) effectively enables the learning of discontinuous values, without requiring manually supervised data as suggested by the current SOTA. More importantly, we show that the close relationship between costates and policies makes the former critical in learning feedback control policies with generalizable safety performance.
@@ 1 Introduction
We consider two-player differential games with general-sum, deterministic dynamics, state constraints, and complete information. The Nash equilibrium values of such games satisfy a set of Hamilton-Jacobi-Isaacs (HJI) equations (Crandall and Lions(1983)). It is well known that approximating values and policies of differential games suffers from the following challenges: Firstly, alleviating the curse of dimensionality (CoD) through physics-informed neural networks (PINN) encounters convergence issues when value discontinuity or large Lipschitz constant is present due to state constraints (Zhang et al.(2023a)Zhang, Ghimire, Zhang, Xu, and Ren): for system states and time starting from which a constraint cannot be satisfied eventually, the value becomes infinite (or large with constraints are treated as penalties). Secondly, for policies to have good safety performance with respect to the state constraints, we need not only small approximation error for values but also for the gradients of values with respect to the states (Yu et al.(2022)Yu, Lu, Meng, and Karniadakis). Lastly, it is often necessary to learn generalizable values and policies across a parametric space of games, e.g., for the inference of game parameters when such information is private (Cardaliaguet(2012)). In this paper we focus on parameters that define large penalties that represent state constraints.
To address these challenges, we propose Pontryagin neural operator (PNO). PNO addresses the first and second challenges by introducing and minimizing costate losses defined on the discrepancy between forward and backward costate and value dynamics following Pontryagin’s Maximum Principle (PMP) (Mangasarian(1966)): First, the costate dynamics that reflect state constraint violation can be readily computed by an ODE solver, and serves as a Lagrangian-frame self-supervised signal to facilitate the learning of highly nonlinear values, which cannot be achieved by the standard Eulerian-frame PINN. Second, the direct connection between costates and policies through the Hamiltonian makes costate-based learning more effective at converging to the ground-truth policies. For the last challenge on generalization to parametric PDEs, we extend our costate-based PINN from solving a single HJI equation to parametric HJIs using a DeepONet architecture (Lu et al.(2021)Lu, Jin, Pang, Zhang, and Karniadakis), a neural operator that supports point-wise value gradient predictions needed for closed-loop control.
@@
(1) Convergence without supervision: The convergence issue in solving discontinuous HJI values via PINN has been investigated in (Zhang et al.(2023a)Zhang, Ghimire, Zhang, Xu, and Ren), where the authors proposed to hybridize PINN using supervised equilibria data generated by solving games offline via PMP with sampled initial states. The major limitation of the hybrid method is in its assumption that knowledge already exists regarding where the informative equilibria should be sampled in the vast state space, e.g., collision and near-collision cases in vehicle interactions. In addition, generating supervisory data via solving PMP encounters numerical and convergence issues due to the existence of multiple equilibria and singular arcs. To the best of the authors’ knowledge, this paper is among the first to address this convergence issue without using any supervised data or domain knowledge of corner cases. To do so, we sample costate trajectories based on a sampling distribution defined over the state space, which evolves along the training according to the costate loss landscape. Considering costate losses as PMP-driven constraints, the sampling distribution essentially captures the dual variables of the PINN. We conduct numerical studies using a two-vehicle game at an uncontrolled intersection. Our experimental results reveal that, with the same data complexity, PNO surpasses the hybrid method in safety performance across a range of parametric collision zones. (2) Solving parametric HJI via value decomposition: This paper is also among the first to study the efficacy of physics-informed neural operators in the context of HJI equations. We show that the DeepONet architecture essentially learns a decomposition of the parametric value and identifies the major value basis functions. Empirical results on the same two-vehicle game show that the safety performance of the learned neural operator across a parametric set of collision zones closely aligns with the ground truth. This investigation opens up new directions towards explainable data-driven models for policy learning: Can we translate basis value (or costate) functions into explainable sub-policies (e.g., temporal logic rules) that together comprise generalizable policies for parametric sets of games?
@@ 2 Related Work
The existence of value for zero-sum differential games with Markovian payoffs and state constraints (and temporal logic specifications in general) have been derived (Fisac et al.(2015)Fisac, Chen, Tomlin, and Sastry). The results have been extended to general-sum games with non-temporal state constraints (Pérolat et al.(2017)Pérolat, Strub, Piot, and Pietquin). To facilitate value approximation using level set methods, an epigraphical technique has been introduced to smooth discontinuous values when state constraints are present (Gammoudi and Zidani(2023)).
@@
HJI equations resulted from differential games are first-order nonlinear PDEs that suffer from CoD when solved by conventional level set methods (Osher et al.(2004)Osher, Fedkiw, and Piechor; Mitchell and Templeton(2005)) or essentially nonoscillatory schemes (Osher and Shu(1991)). Monte Carlo methods, such as variants of PINN (Krishnapriyan et al.(2021)Krishnapriyan, Gholami, Zhe, Kirby, and Mahoney), have recently shown promise at solving high-dimensional PDEs, including HJ equations, provided that the solution is smooth (Weinan et al.(2021)Weinan, Han, and Jentzen). PINN uses a trainable neural net to represent the solution space, and optimizes PDE-driven errors to approximate the solution. Such errors include the boundary residual (Han and Long(2020); Han et al.(2018)Han, Jentzen, and E), the PDE residual (Jagtap et al.(2020)Jagtap, Kawaguchi, and Karniadakis; Bansal and Tomlin(2021)), and supervisory data on the characteristic curves of the PDE (Nakamura-Zimmerer et al.(2021)Nakamura-Zimmerer, Gong, and Kang). Convergence and generalization of PINN have been analyzed under the assumption that both the solution and the network are Lipschitz continuous (Han and Long(2020); Shin et al.(2020)Shin, Darbon, and Karniadakis; Ito et al.(2021)Ito, Reisinger, and Zhang). Recent studies have explored the effectiveness of PINN for solving PDEs with discontinuous solutions (Jagtap et al.(2020)Jagtap, Kawaguchi, and Karniadakis) and solutions with large Kolmogorov n-width (Mojgani et al.(2023)Mojgani, Balajewicz, and Hassanzadeh). However, solving PDEs with discontinuous solutions and only terminal boundaries (such as HJI equations with state constraints) is still an open challenge without prior structural knowledge about the value landscape (Zhang et al.(2023b)Zhang, Ghimire, Zhang, Xu, and Ren).
@@
Neural operators emerged as a promising approach to universally approximating mappings between functions (Kovachki et al.(2023)Kovachki, Li, Liu, Azizzadenesheli, Bhattacharya, Stuart, and Anandkumar), with particularly successful applications to solving parametric PDEs (Wang et al.(2021)Wang, Wang, and Perdikaris). The architecture most related to this paper is DeepONet (Lu et al.(2021)Lu, Jin, Pang, Zhang, and Karniadakis), which is composed of a branch net and a trunk net. The branch net extracts informative features of discretized input functions that define parameters of PDEs and the trunk net captures the basis functions that comprise parametric PDE solutions. DeepONet is one of the neural operator architectures that allow point-wise prediction, instead of predicting the entire function over its input domain (e.g., FNO (Li et al.(2020)Li, Kovachki, Azizzadenesheli, Liu, Bhattacharya, Stuart, and Anandkumar) and GNOT (Hao et al.(2023)Hao, Wang, Su, Ying, Dong, Liu, Cheng, Song, and Zhu)). The extension of neural operators from supervised to PINN training has been studied for solving parametric physics equations defined on 2D and 3D state spaces (Wang et al.(2021)Wang, Wang, and Perdikaris). This paper examines the efficacy of the DeepONet architecture in physics-informed training of values with large Lipschitz constants defined on a 5D state-time space and a 2D parameter space.
3 Differential Games with State Constraints
@@
Let the state dynamics of Player $i$ be $\dot{x}_{i}=f_{i}(x_{i},u_{i})$, where $x_{i}\in\mathcal{X}_{i}\subseteq\mathbb{R}^{d_{x}}$ is the system state and $u_{i}\in\mathcal{U}_{i}\subseteq\mathbb{R}^{d_{u}}$ is the control input. The joint state space is $\mathcal{X}:=\mathcal{X}_{1}\times\mathcal{X}_{2}$. The instantaneous loss of Player $i$ is denoted by $l_{i}(\cdot,\cdot):\mathcal{X}\times\mathcal{U}_{i}\rightarrow\mathbb{R}$ and the terminal loss $g_{i}(\cdot):\mathcal{X}_{i}\rightarrow\mathbb{R}$. The fixed time horizon of the game is $[0,T]$. We use $\textbf{a}_{i}=(a_{i},a_{-i})$ to concatenate elements $a_{i}$ from Player $i$ and $a_{-i}$ from the fellow player, and use $\textbf{a}=(a_{1},a_{2})$. Denote $\alpha_{i}\in\mathcal{A}:\mathcal{X}\times[0,T]\rightarrow\mathcal{U}_{i}$ as Player $i$’s policy. We use $x_{s}^{x_{i},t,\alpha_{i}}$ as the state of Player $i$ at time $s$ if it follows policy $\alpha_{i}$, dynamics $f_{i}$, and starts at $(x_{i},t)$. $\textbf{x}_{s}^{\textbf{x}_{i},t,\boldsymbol{\alpha}_{i}}:=\left(x_{s}^{x_{i},t,\alpha_{i}},x_{s}^{x_{-i},t,\alpha_{-i}}\right)$. Let $c_{i}(\cdot):\mathcal{X}\rightarrow\mathbb{R}$ be a state penalty derived from Player $i$’s state constraints, i.e., for any $\textbf{x}_{i}\in\mathcal{X}$, $c_{i}(\textbf{x}_{i})=0$ if $\textbf{x}_{i}$ satisfies Player $i$’s state constraints, and otherwise $c_{i}(\textbf{x}_{i})$ becomes a large positive number. In this paper, we consider $c_{i}(\cdot)$ to be differentiable but with a large Lipschitz constant. The value of Player $i$ is denoted by $\vartheta_{i}(\cdot,\cdot):\mathcal{X}\times[0,T]\rightarrow\mathbb{R}$. We omit arguments to $f_{i}$, $l_{i}$, $g_{i}$, $c_{i}$, $\vartheta_{i}$ when possible, and decorate them with the superscript $\theta\in\Theta\subseteq\mathbb{R}^{d_{\theta}}$ when the corresponding functions are type-specific, where $\Theta$ is the type space. E.g., $c^{\theta}_{i}(\cdot,\cdot)$ is the state penalty of Player $i$ of type $\theta$. Throughout the paper, we assume that $\mathcal{U}_{i}$ is compact and convex; $f_{i}$ and $c_{i}$ are Lipschitz continuous; $l_{i}$ and $g_{i}$ are Lipschitz continuous and bounded.
@@
The value for Player $i$ is $\vartheta_{i}(\textbf{x}_{i},t)=J_{i}(\textbf{x}_{i},t,\boldsymbol{\alpha}^{\dagger}_{i})$. The HJI equations that govern the values are denoted by $L$, and the boundary conditions by $D$ (Starr and Ho(1969)):
$L(\boldsymbol{\vartheta}_{i},\nabla_{\textbf{x}_{i}}\vartheta_{i},\textbf{x}_{i},t):=\nabla_{t}\vartheta_{i}+\max_{u\in\mathcal{U}_{i}}\left\{\nabla_{\textbf{x}_{i}}\vartheta_{i}^{T}\textbf{f}_{i}-(l_{i}+c_{i})\right\}=0$ (3)
@@
By solving Eq. (3) for $(\vartheta_{1},\vartheta_{2})$, Player $i$’s equilibrium policy can be derived as
$\alpha_{i}^{\dagger}(\textbf{x}_{i},t)=\operatorname*{arg\,max}_{u\in\mathcal{U}_{i}}\left\{\nabla_{\textbf{x}_{i}}\vartheta_{i}^{T}\textbf{f}_{i}-(l_{i}+c_{i})\right\}.$
When needed, we denote by $\mathcal{L}^{\boldsymbol{\theta}}:=(L^{\boldsymbol{\theta}},D^{\boldsymbol{\theta}})$ the HJI of a game parameterized by ${\boldsymbol{\theta}}$.
@@
The characteristic curves of $\mathcal{L}$ are open-loop equilibrium trajectories governed by PMP: For initial state $(\bar{x}_{1},\bar{x}_{2})\in\mathcal{X}$ and $t\in[0,T]$:
@@
Here $\lambda_{i}=\nabla_{x_{i}}\vartheta_{i}$ is the costate of Player $i$. Solving Eq. (4) for a given initial states in $\mathcal{X}\times[0,T]$ is a boundary-value problem (BVP). Provided that convergence can be achieved, the solution to this BVP approximates $\boldsymbol{\vartheta}$ through the resultant characteristic trajectory and the boundary conditions $D$. Zhang et al.(2023a)Zhang, Ghimire, Zhang, Xu, and Ren uses these characteristic trajectories as supervisory data, but assumes that informative initial states, i.e., those for which $c_{i}$ changes significantly along the trajectories, are known. PNO exploits the method of characteristics without requiring prior knowledge. See Sec. 4.

### v2

@@ Abstract
The values of two-player general-sum differential games are viscosity solutions to Hamilton-Jacobi-Isaacs (HJI) equations. Value and policy approximations for such games suffer from the curse of dimensionality (CoD). Alleviating CoD through physics-informed neural networks (PINN) encounters convergence issues when differentiable values with large Lipschitz constants are present due to state constraints. On top of these challenges, it is often necessary to learn generalizable values and policies across a parametric space of games, e.g., for game parameter inference when information is incomplete. To address these challenges, we propose in this paper a Pontryagin-mode neural operator that outperforms the current state-of-the-art hybrid PINN model on safety performance across games with parametric state constraints. Our key contribution is the introduction of a costate loss defined on the discrepancy between forward and backward costate rollouts, which are computationally cheap. We show that the costate dynamics, which can reflect state constraint violation, effectively enables the learning of differentiable values with large Lipschitz constants, without requiring manually supervised data as suggested by the hybrid PINN model. More importantly, we show that the close relationship between costates and policies makes the former critical in learning feedback control policies with generalizable safety performance.
@@ 1 Introduction
We consider two-player general-sum differential games with deterministic dynamics, state constraints, and complete information. The Nash equilibrium values of such games satisfy a set of Hamilton-Jacobi-Isaacs (HJI) equations (Crandall and Lions, 1983; Mitchell et al., 2005; Bressan, 2010). It is well known that approximating values and policies of differential games suffers from the following challenges: Firstly, alleviating the curse of dimensionality (CoD) through physics-informed neural networks (PINN) encounters convergence issues when value discontinuity (or differentiable value with large Lipschitz constant) is present due to state constraints (Zhang et al., 2023b): for system states and time starting from which a constraint cannot be satisfied eventually, the value becomes infinite (or large when constraints are treated as penalties). Secondly, for policies to have good safety performance with respect to the state constraints, we need not only small approximation error for values but also for the gradients of values with respect to the states (Yu et al., 2022). Lastly, it is often necessary to learn generalizable values and policies across a parametric space of games, e.g., for the inference of game parameters when such information is private (Cardaliaguet, 2012). In this paper, we focus on parameters that define large penalties that represent state constraints.
To address these challenges, we propose Pontryagin neural operator (PNO). PNO addresses the first and second challenges by introducing and minimizing costate losses defined on the discrepancy between forward and backward costate and value dynamics following Pontryagin’s Maximum Principle (PMP) (Mangasarian, 1966): First, the costate dynamics that reflect state constraint violation (Bokanowski et al., 2021) can be readily computed by an ODE solver, and serves as a Lagrangian-frame self-supervised signal to facilitate the learning of highly nonlinear values, which cannot be achieved by the standard Eulerian-frame PINN. Second, the direct connection between costates and policies through the Hamiltonian makes costate-based learning more effective at converging to the ground-truth policies. For the last challenge on generalization to parametric PDEs, we extend our costate-based PINN from solving a single HJI equation to parametric HJIs using a DeepONet architecture (Lu et al., 2021), a neural operator that supports point-wise value gradient predictions needed for closed-loop control.
@@
(1) Convergence without supervision: The convergence issue in solving discontinuous HJI values via PINN has been investigated in Zhang et al. (2023a), where the authors proposed to hybridize PINN using supervised equilibria data generated by solving games offline via PMP with sampled initial states. The major limitation of the hybrid method is in its assumption that knowledge already exists regarding where the informative equilibria should be sampled in the vast state space, e.g., collision and near-collision cases in vehicle interactions. In addition, generating supervisory data via solving PMP encounters its own numerical issues due to the existence of multiple equilibria and singular arcs. To the best of the authors’ knowledge, this paper is among the first to address the PINN convergence issue without using any supervised data or domain knowledge of corner cases. To do so, we sample costate trajectories based on a sampling distribution defined over the state space, and evolve this distribution along the training according to the costate loss landscape. Considering costate losses as PMP-driven constraints, the sampling distribution essentially captures the dual variables of the PINN. We conduct numerical studies using a two-vehicle game at an uncontrolled intersection. Our experimental results reveal that, with the same data complexity, PNO surpasses the hybrid neural operator (hybrid PINN + DeepONet) in safety performance across a range of parametric collision zones. (2) Solving parametric HJI via value decomposition: This paper is also among the first to study the efficacy of physics-informed neural operators in the context of parametric HJI equations. We show that the DeepONet architecture essentially learns a decomposition of the parametric value and identifies the major value basis functions. Empirical results on the same two-vehicle game show that the safety performance of the learned neural operator across a parametric set of collision zones closely aligns with the ground truth. This investigation opens up new directions towards explainable data-driven models for policy learning: Can we translate basis value (or costate) functions into explainable sub-policies (e.g., temporal logic rules) that together comprise generalizable policies for parametric sets of games?
@@ 2 Related Work
The existence of value for zero-sum differential games with Markovian payoffs and state constraints (and temporal logic specifications in general) have been derived (Bettiol et al., 2006). The results have been extended to general-sum games with non-temporal state constraints (Zhang et al., 2018). To facilitate value approximation using level set methods, an epigraphical technique has been introduced to smooth discontinuous values when state constraints are present (Gammoudi and Zidani, 2023).
@@
HJI equations resulting from differential games are first-order nonlinear PDEs that suffer from CoD when solved by conventional level set methods (Osher et al., 2004; Mitchell and Templeton, 2005) or essentially nonoscillatory schemes (Osher and Shu, 1991). Monte Carlo methods, such as variants of PINN (Krishnapriyan et al., 2021), have recently shown promise at solving high-dimensional PDEs, including HJ equations, provided that the solution is smooth (E et al., 2021). PINN uses a trainable neural net to represent the solution space, and optimizes PDE-driven errors to approximate the solution. Such errors include the boundary residual (Han and Long, 2020; Han et al., 2018), the PDE residual (Jagtap et al., 2020; Bansal and Tomlin, 2021), and supervisory data on the characteristic curves of the PDE (Nakamura-Zimmerer et al., 2021). Convergence and generalization of PINN have been analyzed under the assumption that both the solution and the network are Lipschitz continuous (Han and Long, 2020; Shin et al., 2020; Ito et al., 2021). Recent studies have explored the effectiveness of PINN for solving PDEs with discontinuous solutions (Jagtap et al., 2020) and solutions with large Kolmogorov n-width (Mojgani et al., 2023). However, solving PDEs with discontinuous solutions and only terminal boundaries (such as HJI equations with state constraints) is still an open challenge without prior structural knowledge about the value landscape (Zhang et al., 2023b).
@@
Neural operators emerged as a promising approach to universally approximating mappings between functions (Kovachki et al., 2023), with particularly successful applications to solving parametric PDEs (Wang et al., 2021). The architecture most related to this paper is DeepONet (Lu et al., 2021), which is composed of a branch net and a trunk net. The branch net extracts informative features of discretized input functions that define parameters of PDEs and the trunk net captures the basis functions that comprise parametric PDE solutions. DeepONet is one of the neural operator architectures that allow point-wise prediction, instead of predicting the entire function over its input domain (e.g., FNO (Li et al., 2020) and GNOT (Hao et al., 2023)). The extension of neural operators from supervised to PINN training has been studied for solving parametric physics equations defined on 2D and 3D state spaces (Wang et al., 2021). This paper examines the efficacy of the DeepONet architecture in physics-informed training of values with large Lipschitz constants defined on a 5D state-time space and a 2D parameter space. We compare two different neural operators empirically: the proposed Pontryagin neural operator and a hybrid one as a baseline.
3 Differential Games with Penalized State Constraints
@@
Let the state dynamics of Player $i$ be $\dot{x}_{i}=f_{i}(x_{i},u_{i})$, where $x_{i}\in\mathcal{X}_{i}\subseteq\mathbb{R}^{d_{x}}$ is the system state and $u_{i}\in\mathcal{U}_{i}\subseteq\mathbb{R}^{d_{u}}$ is the control input. The joint state space is $\mathcal{X}:=\mathcal{X}_{1}\times\mathcal{X}_{2}$. The instantaneous loss of Player $i$ is $l_{i}(\cdot,\cdot):\mathcal{X}\times\mathcal{U}_{i}\rightarrow\mathbb{R}$ and the terminal loss $g_{i}(\cdot):\mathcal{X}_{i}\rightarrow\mathbb{R}$. The fixed time horizon of the game is $[0,T]$. We use $\textbf{a}_{i}=(a_{i},a_{-i})$ to concatenate elements $a_{i}$ from Player $i$ and $a_{-i}$ from the fellow player, and use $\textbf{a}=(a_{1},a_{2})$. Denote $\alpha_{i}\in\mathcal{A}:\mathcal{X}\times[0,T]\rightarrow\mathcal{U}_{i}$ as Player $i$’s policy. We use $x_{s}^{x_{i},t,\alpha_{i}}$ as the state of Player $i$ at time $s$ if he follows policy $\alpha_{i}$, dynamics $f_{i}$, and starts at $(x_{i},t)$. $\textbf{x}_{s}^{\textbf{x}_{i},t,\boldsymbol{\alpha}_{i}}:=\left(x_{s}^{x_{i},t,\alpha_{i}},x_{s}^{x_{-i},t,\alpha_{-i}}\right)$. Let $c_{i}(\cdot):\mathcal{X}\rightarrow\mathbb{R}$ be a state penalty derived from Player $i$’s state constraints, i.e., for any $\textbf{x}_{i}\in\mathcal{X}$, $c_{i}(\textbf{x}_{i})=0$ if $\textbf{x}_{i}$ satisfies Player $i$’s state constraints, and otherwise $c_{i}(\textbf{x}_{i})$ becomes a large positive number. In this paper, we consider $c_{i}(\cdot)$ to be differentiable but with a large Lipschitz constant. The value of Player $i$ is denoted by $\vartheta_{i}(\cdot,\cdot):\mathcal{X}\times[0,T]\rightarrow\mathbb{R}$. We omit arguments to $f_{i}$, $l_{i}$, $g_{i}$, $c_{i}$, $\vartheta_{i}$ when possible, and decorate them with the superscript $\theta\in\Theta\subseteq\mathbb{R}^{d_{\theta}}$ when the corresponding functions are type-specific, where $\Theta$ is the type space. E.g., $c^{\theta}_{i}(\cdot,\cdot)$ is the state penalty of Player $i$ of type $\theta$. Throughout the paper, we assume that $\mathcal{U}_{i}$ is compact and convex; $f_{i}$ and $c_{i}$ are Lipschitz continuous; $l_{i}$ and $g_{i}$ are Lipschitz continuous and bounded.
@@
The value for Player $i$ is $\vartheta_{i}(\textbf{x}_{i},t)=J_{i}(\textbf{x}_{i},t,\boldsymbol{\alpha}^{\dagger}_{i})$. The HJI equations that govern the values, denoted by $L$, and the boundary conditions, by $D$, are the following Starr and Ho (1969):
$L(\vartheta_{i},\nabla_{\textbf{x}_{i}}\vartheta_{i},\textbf{x}_{i},t):=\nabla_{t}\vartheta_{i}+\max_{u\in\mathcal{U}_{i}}\left\{\nabla_{\textbf{x}_{i}}\vartheta_{i}^{T}\textbf{f}_{i}-(l_{i}+c_{i})\right\}=0$ (3)
@@
Therefore, Player $i$’s equilibrium policy can be derived as $\alpha_{i}^{\dagger}(\textbf{x}_{i},t)=\operatorname*{arg\,max}_{u\in\mathcal{U}_{i}}\{\nabla_{\textbf{x}_{i}}\vartheta_{i}^{T}\textbf{f}_{i}-(l_{i}+c_{i})\}$ if Eq. (3) can be solved for $(\vartheta_{1},\vartheta_{2})$ (Bressan, 2010). When needed, we denote by $\mathcal{L}^{\boldsymbol{\theta}}:=(L^{\boldsymbol{\theta}},D^{\boldsymbol{\theta}})$ the HJI of a game parameterized by ${\boldsymbol{\theta}}$.
@@
The characteristic curves of $\mathcal{L}$ are open-loop equilibrium trajectories governed by PMP, which entails the following for initial state $(\bar{x}_{1},\bar{x}_{2})\in\mathcal{X}$ and $t\in[0,T]$:
@@
Here $\lambda_{i}=\nabla_{x_{i}}\vartheta_{i}$ is the costate of Player $i$. Solving Eq. (4) for a given initial states in $\mathcal{X}\times[0,T]$ is a boundary-value problem (BVP). Provided that convergence can be achieved, the solution to this BVP approximates $\boldsymbol{\vartheta}$ through the resultant characteristic trajectory and the boundary conditions $D$. Zhang et al. (2023a) uses these characteristic trajectories as supervisory data, but assumes that informative initial states, i.e., those for which $c_{i}$ changes significantly along the trajectories, are known. PNO exploits the method of characteristics without requiring prior knowledge. See Sec. 4.